# Personalized Negative Reservoir for Incremental Learning in Recommender Systems

### v1

@@ Abstract.
Recommender systems have become an integral part of online platforms. Every day the volume of training data is expanding and the number of user interactions is constantly increasing. The exploration of larger and more expressive models has become a necessary pursuit to improve user experience. However, this progression carries with it an increased computational burden. In commercial settings, once a recommendation system model has been trained and deployed it typically needs to be updated frequently as new client data arrive. Cumulatively, the mounting volume of data is guaranteed to eventually make full batch retraining of the model from scratch computationally infeasible. Naively fine-tuning solely on the new data runs into the well-documented problem of catastrophic forgetting. Despite the fact that negative sampling is a crucial part of training with implicit feedback, no specialized technique exists that is tailored to the incremental learning framework. In this work, we take the first step to propose, a personalized negative reservoir strategy which is used to obtain negative samples for the standard triplet loss. This technique balances alleviation of forgetting with plasticity by encouraging the model to remember stable user preferences and selectively forget when user interests change. We derive the mathematical formulation of a negative sampler to populate and update the reservoir. We integrate our design in three SOTA and commonly used incremental recommendation models. We show that these concrete realizations of our negative reservoir framework achieve state-of-the-art results in standard benchmarks, on multiple standard top-k evaluation metrics.
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title; June 03–05, 2018; Woodstock, NY
Figure 1. Toy example with daily model updates with 3 item categories: pet items, car part items and coffee items. The figure shows a user’s interactions with these categories over a week. (a) User $u$ interacts with more items from pet item category in the start of the start of the week but his interests change over time and he interacts with more car part items by the end of the week. (b) The change of preference is reflected on the histograms of user-item categories $\mathbf{H}_{u,t}$. (c) The histograms are normalized $\widehat{\mathbf{H}}_{u,t}$ and projected on the simplex $\Delta_{2}$. There we see the interest shift on the simplex representing the user’s interest shift from $t=2,$ (Tuesday) to $t=3$, (Wednesday) and on the right the user’s interest shift from $t=3,$ (Wednesday) to $t=4$, (Thursday). We see that the user exhibits a large change in interests on the right (moves far away from current neighborhood in simplex denoted by a red dashed circle). (d) We propose to sample more negatives from pet category when fine tuning the model on Thursday to allow the model to quickly adjust to new user interests (in the figure the right bucket has many more pet related items).
1. Introduction
@@
Recommender systems have become a crucial part of online services. Delivering highly relevant item recommendations not only enhances user experience but also bolsters the revenue of service providers. The advent of deep learning-based recommender systems (Covington et al., 2016; Guo et al., 2017; Cheng et al., 2016; He et al., 2023) has significantly elevated the quality of user and item representations. To more accurately represent user behavior, there has been a substantial expansion in the volume of training data, accumulated from the long user-item interaction history (He et al., 2023; Xu et al., 2020). Thus, the exploration of larger and more expressive models has become a vital research direction. For example, Graph Neural Network (GNN) (Kipf and Welling, 2017) based recommendation methods can achieve compelling performance on recommendation tasks because of their ability to model the rich relational information of the data through the message passing paradigm (van den Berg et al., 2018; Wang et al., 2019a, 2021a; Sun et al., 2019). However, this evolution brings with it a potential increase in computational burden. An industrial-scale recommendation serving model, once integrated into an online system, usually requires regular updates to accommodate the arrival of recent client data. The constant arrival of new data inevitably leads to a point where full-batch retraining of the model from scratch becomes infeasible.
@@
One straightforward way to tackle this computational challenge is to train the backbone model in an incremental fashion, updating it only when a new data block arrives, instead of full batch retraining with older data. In industrial-level recommender systems, the new data block can arrive in a daily, hourly or even a shorter interval (Xu et al., 2020; Wang et al., 2021b), depending on the application. Unfortunately, naively fine-tuning the model with new data leads to the well-known issue of “catastrophic forgetting” (Kirkpatrick et al., 2017), i.e., the model discards information from earlier data blocks and overfits to the newly acquired data. There are two mainstream methods to alleviate the catastrophic forgetting problem: (i) experience (reservoir) replay (Prabhu et al., 2020; Ahrabian et al., 2021), and (ii) regularization-based knowledge distillation (Castro et al., 2018; Kirkpatrick et al., 2017; Xu et al., 2020; Wang et al., 2021b, 2023). Reservoir replay methods retrain on some previously observed user interactions from past data blocks while jointly training with the new data. The regularization techniques are typically formulated as a knowledge distillation problem where the model trained on the old data takes the role of the “teacher” model and the model fine-tuned on the new data is regarded as the “student” model. A knowledge distillation loss is applied to preserve the information from the teacher to the student model through model weight distillation (Castro et al., 2018; Kirkpatrick et al., 2017), structural (Xu et al., 2020), or contrastive distillation (Wang et al., 2021b).
@@
However, there is one important aspect of the incremental learning framework that has received very little attention. Negative sampling plays a critical role in recommendation system training. The most common strategy for negative sampling involves uniform random sampling of the negatives (Rendle et al., 2009). While subsequent works improve upon the negative sampler for the static setting (one-time learning), there are no works dedicated to addressing negative sampling in the incremental learning setting. We identify two unique challenges for designing a good negative sampler for an incremental learning framework. First, the negative reservoir must be personalized for each user, and notably, it should model a user’s interests or preference shift across consecutive time blocks. This can provide a significant distribution bias in the negative sampling process. This relates to a classic trade-off in online learning, achieving a correct balance in retaining knowledge learned from prior training (stability), and being flexible enough to learn concepts from new observations (plasticity). Second, the ranking list (prediction) produced by the model from the previous time block should be exploited in order to find informative negative samples for the training in the new block.
@@
Considering the above-mentioned challenges, we design the first negative reservoir strategy tailored for the incremental learning framework. This negative reservoir contains the most effective negative samples in each incremental training block based on the user change of interest. Our negative reservoir design is compatible with almost all existing incremental learning frameworks for recommender systems. In our experiments, we integrate our negative reservoir design into three recent incremental learning frameworks. Our designed negative reservoir achieves state-of-the-art performance when incorporated in three standard incremental learning frameworks, improving GraphSAIL (Xu et al., 2020), SGCT (Wang et al., 2021b), and LWC-KD (Wang et al., 2021b) by an average of 13.4%, 9.4% and 6.7%, respectively, across five large-scale recommender system datasets. Our main contributions can be summarized as:
@@
  • This is the first work to propose a negative reservoir design tailored for incremental learning in recommender systems. The approach, for which we provide a principled mathematical derivation, can be easily incorporated into almost all existing incremental learning frameworks that involve triplet loss.
  • We propose a personalized negative reservoir based on the user-specific preference change. All prior works assume static user preferences when selecting negative items.
@@
  • We demonstrate the effectiveness of our personalized negative reservoir via a thorough comparison on five diverse datasets. We strongly and consistently improve upon recent SOTA incremental learning techniques using our negative reservoir framework. These results are not achievable with other negative samplers.
2. Related Work
2.1. Incremental Learning for Rec. Systems
@@ 2. Related Work
Incremental learning is a training strategy that can allow models to update continually when new data arrives. However, naively fine-tuning the model with new data leads to “catastrophic forgetting” (Kirkpatrick et al., 2017; Shmelkov et al., 2017; Castro et al., 2018), i.e., the model overfits to the newly acquired data and loses the ability to generalize well on data from previous blocks. There are two main research branches that aim to alleviate the catastrophic forgetting issue. The first direction is called reservoir replay or experience replay. Well-acknowledged works such as iCarl (Rebuffi et al., 2017) and GDumb (Prabhu et al., 2020) construct a reservoir from the old data and replay the reservoir while training with the new data. The reservoir is usually constructed via direct optimization or greedy heuristics. A recent incremental framework for graph-based recommender systems extended the core idea of the GDumb heuristic and proposed a reservoir sampler based on the node degree (Ahrabian et al., 2021). Another line of research focuses on regularization-based knowledge distillation. The model trained using old data blocks serves as the teacher model and the model fine-tuned using the new data is the student. A knowledge distillation (KD) loss (Hinton et al., 2015) is applied to preserve certain properties that were learned from the historical data. In GraphSAIL (Xu et al., 2020), each node’s local and global structure in the user-item bipartite graph is preserved. By contrast, in SGCT and LWC-KD (Wang et al., 2021b), a layer-wise contrastive distillation loss is applied to enable intermediate layer embeddings and structure-aware knowledge to be transferred successfully between the teacher model and the student model. However, one important aspect of the incremental learning framework that has received very little attention is how to properly design a negative sample reservoir. This is a key omission considering the important role of negative sampling in recommendation. In this paper, we shed some light on how to design a negative reservoir that is specifically tailored to the special characteristics of incremental learning in recommender systems.
2.2. Negative Sampling in Rec. Systems
@@
Since the number of non-observed interactions in a recommendation dataset is vast (often in the billions (Ying et al., 2018a)), sampling a small number of negative items is necessary for efficient learning. Random negative sampling is the default sampling strategy in Bayesian Personalized Ranking (BPR) (Rendle et al., 2009). Some more recent attempts (Rendle and Freudenthaler, 2014a; Caselles-Dupré et al., 2018; Zhao et al., 2015; Ying et al., 2018a) aim to design a better heuristic negative sampling strategy to obtain more effective negative samples from non-interacting items. In general, the intuition is that presenting “harder” negative samples during training should encourage the model to learn better item and user representations. Some heuristics select negative samples based on the popularity of the item (Rendle and Freudenthaler, 2014a) or the node degree (Caselles-Dupré et al., 2018). Some strive to identify hard negative samples by rejection (Zhao et al., 2015), or via personalized PageRank (Ying et al., 2018a). Other works focus on a more sophisticated model-based negative sampler. For example, DNS (Zhang et al., 2013) chooses negative samples from the top ranking list produced by the current prediction model. IRGAN (Wang et al., 2017) uses a minimax game realized by a generative adversarial network framework to produce negative sample candidates. Yu et al. (Yu et al., 2022) address the issue of class imbalance of negatives in a static recommendation setting. Although these negative sampling strategies yield improvement when applied naively to incremental recommendation, they do not take user time evolving interests into account and as such leave substantial room for improvement in this specific setting. Additionally, the GAN techniques often rely on reinforcement learning for the optimization. In industrial settings the instability of GAN optimization can make it challenging to use for real world applications.
3. Problem Statement
@@
In this section, we provide a clear definition of the incremental learning setting and our problem statement. Consider a bipartite graph $\mathcal{G}_{t}=(\mathcal{U}_{t},\mathcal{I}_{t},\mathcal{E}_{t})$ with a node set $\mathcal{V}_{t}=\mathcal{U}_{t}\bigcup\mathcal{I}_{t}$ that consists of two types: user nodes $\mathcal{U}_{t}$ and item nodes $\mathcal{I}_{t}$. A set of edges $\mathcal{E}_{t}$ interconnects elements of $\mathcal{U}_{t}$ and $\mathcal{I}_{t}$; thus each edge of $\mathcal{G}_{t}$ encodes one user-item interaction. We consider learning in discrete intervals, and use the integer $t$ to index the $t$-th interval. This corresponds to a continuous time interval $[t\Delta T,(t+1)\Delta T)$. When we refer to the interactions at time t, indicated as $\mathcal{E}_{(t,t+1]}$, we thus mean all interactions in the interval $[t\Delta T,(t+1)\Delta T)$. The graph and its component nodes and edges are indexed by integer time $t$ as they evolve over time in a discrete fashion. The graph structure is updated in regular intervals according to the following rule:
(1) $\mathcal{G}_{t+1}=\left(\mathcal{U}_{t}\cup\mathcal{U}_{(t,t+1]},\mathcal{I}_{t}\cup\mathcal{I}_{(t,t+1]},\mathcal{E}_{(t,t+1]}\right),$
@@
where $\mathcal{U}_{t},\mathcal{I}_{t},\mathcal{E}_{t}$ represent the cumulative user-item interactions up to and including time $t\Delta T$ and $\mathcal{U}_{(t,t+1]},\mathcal{I}_{(t,t+1]},\mathcal{E}_{(t,t+1]}$ represent the user-item interactions accrued during the time interval $[t\Delta T,(t+1)\Delta T)$. Note that in our setting we do not construct a graph using the old set of edges $\mathcal{E}_{t}$; we merely have the observed edges from the current time block and all the user and item nodes from previous and current time blocks. For the rest of this analysis, the user-item interactions in the interval $(0,t]$ are called “base block data” and interactions belonging to subsequent time intervals $\{(t,t+1],(t+1,t+2],\dots\}$ are referred to as “incremental block data”. The goal resembles a standard recommendation task; given $\mathcal{G}_{t}=(\mathcal{U}_{t},\mathcal{I}_{t},\mathcal{E}_{(t-1,t]})$, we are expected to provide a matrix $\mathbf{R}^{\lvert\mathcal{U}_{t}\rvert\times\lvert\mathcal{I}_{t}\rvert}$ ranking all items in order of relevance to each user. The main difference with respect to a standard static recommendation task is that, due to the temporal nature of the incremental learning, our training set includes data up to block $t_{T}$ but we aim to predict item ranking for time $t_{T+\Delta T}$. To measure the quality of recommended items we employ the standard evaluation metrics for the “topK” recommendation task, Recall@K, Precision@K, MAP@K, NDCG@K.
@@
We note the difference between our setting and the distinct problem of sequential learning. Incremental learning aims to ameliorate the computational bottleneck for training, which usually limits the training instances to the most recent time block. To better inherit knowledge from the past data and the previously trained model, a specially designed knowledge distillation or experience (reservoir) replay is usually applied. In contrast, the sequential recommendation problem focuses on designing time-sensitive encoders (Hidasi and Karatzoglou, 2018; Kang and McAuley, 2018; Fan et al., 2021) (e.g., memory units, attention mechanisms) to better capture users’ short and long-term preferences. Thus, incremental learning is a training strategy while sequential learning focuses on specific model design for sequential data. Indeed, incremental learning training strategies can be applied to different types of backbone models that may be sequential or non-sequential.
4. Preliminaries

### v2

@@ Abstract
Recommender systems have become an integral part of online platforms. Every day the volume of training data is expanding and the number of user interactions is constantly increasing. The exploration of larger and more expressive models has become a necessary pursuit to improve user experience. However, this progression carries with it an increased computational burden. In commercial settings, once a recommendation system model has been trained and deployed it typically needs to be updated frequently as new client data arrive. Cumulatively, the mounting volume of data is guaranteed to eventually make full batch retraining of the model from scratch computationally infeasible. Naively fine-tuning solely on the new data runs into the well-documented problem of catastrophic forgetting. Despite the fact that negative sampling is a crucial part of training with implicit feedback, no specialized technique exists that is tailored to the incremental learning framework. In this work, we propose a personalized negative reservoir strategy, which is used to obtain negative samples for the standard triplet loss of graph-based recommendation systems. Our technique balances alleviation of forgetting with plasticity by encouraging the model to remember stable user preferences and selectively forget when user interests change. We derive the mathematical formulation of a negative sampler to populate and update the reservoir. We integrate our design in three SOTA and commonly used incremental recommendation models. We show that these concrete realizations of our negative reservoir framework achieve state-of-the-art results for standard benchmarks using multiple top-k evaluation metrics.
1 Introduction
@@
Recommender systems have become a crucial part of online services. Delivering highly relevant item recommendations not only enhances user experience but also bolsters the revenue of service providers. The advent of deep learning-based recommender systems has significantly elevated the quality of user and item representations (Covington et al., 2016; Guo et al., 2017; Cheng et al., 2016; He et al., 2023). To more accurately represent user behavior, there has been a substantial expansion in the volume of training data, accumulated from the long user-item interaction history (He et al., 2023; Xu et al., 2020). Thus, the exploration of larger and more expressive models has become a vital research direction. For example, Graph Neural Network (GNN) based recommendation methods can achieve compelling performance on recommendation tasks because of their ability to model the rich relational information of the data through the message passing paradigm (van den Berg et al., 2018; Wang et al., 2019b; 2021a; Sun et al., 2019).
@@
However, this evolution brings with it a potential increase in computational burden. An industrial-scale recommendation serving model, once integrated into an online system, usually requires regular updates to accommodate the arrival of recent client data. The constant arrival of new data inevitably leads to a point where full-batch retraining of the model from scratch becomes infeasible. One straightforward way to tackle this computational challenge is to train the backbone model in an incremental fashion, updating it only when a new data block arrives, instead of full batch retraining with older data. In industrial-level recommender systems, the new data block can arrive in a daily, hourly or even in a shorter interval (Xu et al., 2020; Wang et al., 2021b), depending on the application. Unfortunately, naively fine-tuning the model with new data leads to the well-known issue of “catastrophic forgetting” (Kirkpatrick et al., 2017), i.e., the model discards information from earlier data blocks and overfits to the newly acquired data. There are two mainstream methods to alleviate the catastrophic forgetting problem: (i) experience (reservoir) replay (Prabhu et al., 2020; Ahrabian et al., 2021); and (ii) regularization-based knowledge distillation (Castro et al., 2018; Kirkpatrick et al., 2017; Xu et al., 2020; Wang et al., 2021b; 2023). Reservoir replay methods retrain on some previously observed user interactions from past data blocks while jointly training with the new data. Regularization techniques are typically formulated as a knowledge distillation problem where the model trained on the old data takes the role of the “teacher” model and the model fine-tuned on the new data is regarded as the “student” model. A knowledge distillation loss is applied to preserve the information from the teacher to the student model through model weight (Castro et al., 2018; Kirkpatrick et al., 2017), structural (Xu et al., 2020), or contrastive distillation (Wang et al., 2021b).
@@
We note there is one important aspect of the incremental learning framework for graph based recommender systems that has received very little attention. Negative sampling plays a critical role in recommendation system training. This is because when training with implicit feedback we do not have access to explicit user dislikes so we need a mechanism to select low-user interest items for training. Sampling negative items is necessary as training on all possible positive and negative item pairs would be too costly. A good negative sampler can lead to increased convergence speed. This is key in live deployment settings where there is limited amount of time to train before a new batch of data is generated. The most common strategy for negative sampling involves uniform sampling (Rendle et al., 2009). While subsequent works improve upon the negative sampler for the static setting, there are no works dedicated to addressing negative sampling in incremental learning. We identify two unique challenges for designing a good negative sampler for an incremental learning framework. First, the negative reservoir must be personalized for each user and it should model a user’s interests or preference shift across time blocks. This can provide a significant distribution bias in the negative sampling process. This relates to a classic trade-off in continual learning, namely achieving a correct balance between retaining knowledge learned from prior training (stability), and being flexible enough to adapt to new patterns (plasticity) in a delta-regime (Oreshkin et al., 2024) that combines the two. Second, the ranking (prediction) produced by the model from the previous time block should be exploited as a baseline for informative negative samples in the new block. We note that the focus of our work is on accounting for user interest shifts rather addressing the problem from a traditional continual learning perspective that focuses on preventing catastrophic forgetting (Wang et al., 2023). The reason for this is that, as we show in the experiments and case study, existing incremental learning techniques are too focused on stability and neglect model plasticity.
@@
Considering the above-mentioned challenges, we design the first negative reservoir strategy tailored for the incremental learning framework. This negative reservoir contains the most effective negative samples in each incremental training block based on the user change of interest. Methodologically this translates into two hypotheses concerning the properties of a good sampling distribution for negative items. Firstly, the negative sampler should respond to user changes of interest and sample more items from item categories that the user is losing interest in over time and secondly the sampler should prefer items that are ranked highly (hard negatives) to induce larger gradients and learn faster. Our negative reservoir design is compatible with any incremental learning framework that employs negative sampling, which includes the standard incremental learning frameworks for recommender systems. In our experiments, we integrate our negative reservoir design into three recent incremental learning frameworks. Our designed negative reservoir achieves state-of-the-art performance when incorporated in three standard incremental learning frameworks, improving GraphSAIL (Xu et al., 2020), SGCT (Wang et al., 2021b), and LWC-KD (Wang et al., 2021b) by an average of 11.2%, 8.3% and 6.4%, respectively, across six large-scale recommender system datasets. Our main contributions can be summarized as:
@@
  • This is the first work to propose a negative reservoir design tailored for incremental learning in graph-based recommender systems. The approach, for which we provide a principled mathematical derivation in Section 5.1, is compatible with existing learning frameworks that involve triplet loss.
@@
  • We demonstrate the effectiveness of our personalized negative reservoir via a thorough comparison on six diverse datasets. We strongly and consistently improve upon recent SOTA incremental learning techniques (Sections 6.3). These results are not achievable with other negative samplers (Section 6.4).
  • We propose a personalized negative reservoir based on the user-specific preference change. All prior works assume static user interests when selecting negative items. In Section 6.6 we demonstrate that for users with high interest shift our model can quickly adapt and clearly improves recommendations.
2 Related Work
2.1 Incremental Learning for Recommendation Systems
Incremental learning is a training strategy that allows models to update continually as new data arrive. However, naively fine-tuning the model with new data leads to “catastrophic forgetting” (Kirkpatrick et al., 2017; Shmelkov et al., 2017; Castro et al., 2018), i.e., the model overfits to the newly acquired data and loses the ability to generalize well on data from previous blocks. There are two main research branches that aim to alleviate the catastrophic forgetting issue. The first direction is called reservoir replay or experience replay. Well known works such as iCarl (Rebuffi et al., 2017) and GDumb (Prabhu et al., 2020) construct a reservoir from the old data and replay the reservoir while training with the new data. The reservoir is usually constructed via direct optimization or heuristics. In the next paragraph we discuss methods related to incremental learning for graph-based recommendation systems. We note that there is extensive literature outside this area, for example in online learning for CTR prediction (Aharon et al., 2017) and general online or incremental learning (Valkanas et al., 2024; Waxman & Djuric, 2024) that we do not cover because it is beyond the scope of this work.
@@ 2.1 Incremental Learning for Recommendation Systems
A recent incremental framework for graph-based recommender systems extended the core idea of the GDumb heuristic and proposed a reservoir sampler based on node degrees (Ahrabian et al., 2021). Another line of research focuses on regularization-based knowledge distillation (KD). The model trained using old data blocks serves as the teacher model, and the model that is fine-tuned using the new data is the student. A KD loss (Hinton et al., 2015) is applied to preserve certain properties that were learned from the historical data. In GraphSAIL (Xu et al., 2020), each node’s local and global structure in the user-item bipartite graph is preserved. By contrast, in SGCT and LWC-KD (Wang et al., 2021b), a layer-wise contrastive distillation loss is applied to enable intermediate layer embeddings and structure-aware knowledge to be transferred between the teacher model and the student model. However, one important aspect of the incremental learning framework that has received very little attention is how to properly design a negative sample reservoir. This is a key omission considering the important role of negative sampling in recommendation. In this paper, we shed some light on how to design a negative reservoir tailored to the special characteristics of incremental learning in recommender systems.
2.2 Negative Sampling in Recommendation Systems
@@
Since the number of non-observed interactions in a recommendation dataset is vast (often in the billions (Ying et al., 2018a)), sampling a small number of negative items is necessary for efficient learning. Random negative sampling is the default sampling strategy in Bayesian Personalized Ranking (BPR) (Rendle et al., 2009). Some more recent attempts aim to design a better heuristic negative sampling strategy to obtain more effective negative samples from non-interacting items (Rendle & Freudenthaler, 2014; Caselles-Dupré et al., 2018; Zhao et al., 2015; Ying et al., 2018a). In general, the intuition is that presenting “harder” negative samples during training should encourage the model to learn better item and user representations. Some heuristics select negative samples based on the popularity of the item (Rendle & Freudenthaler, 2014) or the node degree (Caselles-Dupré et al., 2018). Some strive to identify hard negative samples by rejection (Zhao et al., 2015), or via personalized PageRank (Ying et al., 2018a). Other works focus on a more sophisticated model-based negative sampler. For example, DNS (Zhang et al., 2013) chooses negative samples from the top ranking list produced by the current prediction model. IRGAN (Wang et al., 2017) uses a minimax game realized by a generative adversarial network framework to produce negative sample candidates. Yu et al. (2022) address the issue of class imbalance of negatives in a static recommendation setting. While not strictly proposed for recommendation systems, Yao et al. (2022) and Chen et al. (2023) propose negative samplers for knowledge graphs that aim to sample “hard” negatives while minimizing the chance of sampling false negatives, i.e., unobserved true positives. Although these negative sampling strategies yield improvement when applied naively to incremental recommendation, they do not take the time-evolving interests of users into account and as such leave substantial room for improvement in this specific setting. Additionally, the GAN techniques often rely on reinforcement learning for the optimization. In industrial settings, the instability of GAN optimization is highly undesirable.
3 Problem Statement
@@
In this section, we provide a clear definition of the incremental learning setting and our problem statement. Consider a bipartite graph $\mathcal{G}_{t}=(\mathcal{U}_{t},\mathcal{I}_{t},\mathcal{E}_{t})$ with a node set $\mathcal{V}_{t}=\mathcal{U}_{t}\bigcup\mathcal{I}_{t}$ that consists of two types: user nodes $\mathcal{U}_{t}$ and item nodes $\mathcal{I}_{t}$. A set of edges $\mathcal{E}_{t}$ interconnects elements of $\mathcal{U}_{t}$ and $\mathcal{I}_{t}$; thus each edge of $\mathcal{G}_{t}$ encodes one user-item interaction. As is done frequently, in a recommendation setting, we only have access to implicit feedback data. Concretely, this means that for each user $u$ we have access to the set of items that a user interacts with $\mathcal{I}_{u}^{+}=\{i:(u,i)\in\mathcal{E}\}$ but no explicit set of user dislikes $\mathcal{I}_{u}^{-}$. We consider learning in discrete intervals, and use the integer $t$ to index the $t$-th interval. This corresponds to a continuous time interval $[t\Delta T,(t+1)\Delta T)$. When we refer to the interactions at time t, indicated as $\mathcal{E}_{(t,t+1]}$, we thus mean all interactions in the interval $[t\Delta T,(t+1)\Delta T)$. The graph and its component nodes and edges are indexed by integer time $t$ as they evolve over time in a discrete fashion. The graph update rule is:
$\mathcal{G}_{t+1}=\left(\mathcal{U}_{t}\cup\mathcal{U}_{(t,t+1]},\mathcal{I}_{t}\cup\mathcal{I}_{(t,t+1]},\mathcal{E}_{(t,t+1]}\right),$ (1)
@@
where $\mathcal{U}_{t},\mathcal{I}_{t},\mathcal{E}_{t}$ represent the cumulative user-item interactions up to and including time $t\Delta T$ and $\mathcal{U}_{(t,t+1]},\mathcal{I}_{(t,t+1]},\mathcal{E}_{(t,t+1]}$ represent the user-item interactions accrued during the time interval $[t\Delta T,(t+1)\Delta T)$. Note that in our setting we do not construct a graph using the old set of edges $\mathcal{E}_{t}$; we merely have the observed edges from the current time block and all the user and item nodes from previous and current time blocks. For the rest of this analysis, the user-item interactions in the interval $(0,t]$ are called “base block data” and interactions belonging to subsequent time intervals $\{(t,t+1],(t+1,t+2],\dots\}$ are referred to as “incremental block data”. The goal resembles a standard recommendation task; given $\mathcal{G}_{t}=(\mathcal{U}_{t},\mathcal{I}_{t},\mathcal{E}_{(t-1,t]})$, we are expected to provide a matrix $\mathbf{R}^{\lvert\mathcal{U}_{t}\rvert\times\lvert\mathcal{I}_{t}\rvert}$ ranking all items in order of relevance to each user. The main difference with respect to a standard static recommendation task is that, due to the temporal nature of the incremental learning, our training set includes data up to block $t_{T}$ but we aim to predict item ranking for time $t_{T+\Delta T}$. To measure the quality of recommended items we employ the standard evaluation metrics for the “topK” recommendation task, Recall@K, Precision@K, MAP@K, NDCG@K (definitions in Appendix A).
@@
We note the difference between our setting and the distinct problem of sequential learning. Incremental learning aims to ameliorate the computational bottleneck for training, which usually limits the training instances to the most recent time block. To better inherit knowledge from the past data and the previously trained model, a specially designed knowledge distillation or experience (reservoir) replay is usually applied. In contrast, the sequential recommendation problem focuses on designing time-sensitive encoders (Hidasi & Karatzoglou, 2018; Kang & McAuley, 2018; Fan et al., 2021) (e.g., memory units, attention mechanisms) to better capture users’ short and long-term preferences. Thus, incremental learning is a training strategy, whereas sequential learning focuses on specific model design for sequential data. Indeed, incremental learning training strategies can be applied to different types of backbone models that may be sequential or not. Our negative reservoir tracks shifts in user interest over time; however, we do not apply this approach to new users, as there is no historical interaction data available to observe such shifts. We consider the cold-start problem to be distinct from our setting. We note that, since our framework assumes small and frequent updates for each data block, the proportion of new users and items within each incremental training block is generally small. The cold-start user and item representations are learned via the standard GNN backbone and improving performance for them is beyond our scope.
4 Preliminaries